Company A leased equipment it purchased for $600,000 to Company B. The arrangement specifies annual payments beginning at the20
Answer #1

Ans:

Machine costs : $600,000

Buyback value after 4 years : $120,000

PV value @8% for 4 years : 1/(1.08)^4 = 0.73503

PV of buyback value : $120,000 * 0.73503 = $88,203

Present Value of total payments : $600,000 - $88,203 = $511,796

Annual Lease Payment : PV of Total payments / Annuity value @8% for 4 years

Annuity value @8% for 4 years : 1/(1.08)^1 + 1/(1.08)^2 + 1/(1.08)^3 + 1/(1.08)^4 = 3.31217

Annual Lease Payment : $511,796 / 3.31217 = $154,522

So correct answer option C.

You can ask your queries in comment section. Please rate if you found this helpful.